OpenStudy (anonymous):

What is the simplest form of the radical expression? 4^(3)√(3x)+ 5^(3) √(10x) Im getting 64√3x+125√10x

OpenStudy (anonymous):

but thats not an answer choice

OpenStudy (anonymous):

@hartnn @mathlover2014 @ranga @Embryo @jcpd910

OpenStudy (anonymous):

4^3√3x + 5^3√10x = x(4^3√3 + 5^3√10)

OpenStudy (anonymous):

Could you use the equation button to create the equation for us please? :)

hartnn (hartnn):

what are your answer choices ?

OpenStudy (anonymous):

who?

hartnn (hartnn):

jane, you factored out \(\sqrt x\), right ?

OpenStudy (anonymous):

\[4^{3}\sqrt{3x}+5^{3}\sqrt{10x}\]

hartnn (hartnn):

post the answer choices please :)

hartnn (hartnn):

FIRSTLY they are cube roots! \(\Large \sqrt[3]{3x}\) Secondly, 2 cube roots cannot add up, and get simplified...

OpenStudy (anonymous):

so D!

hartnn (hartnn):

correct! :D

OpenStudy (anonymous):

thank you, i was overthinking it lol

hartnn (hartnn):

no problem,it happens :) welcome ^_^

OpenStudy (anonymous):

Would you help with one more please? @hartnn

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

\[\sqrt[4]{81} x ^{20} y ^{8}\]

OpenStudy (anonymous):

What is a simpler form of the radical expression? @hartnn

hartnn (hartnn):

that forth root is entire right ?? covering x and y too ?

hartnn (hartnn):

\((\sqrt[4]{81 x ^{20} y ^{8}}\))

hartnn (hartnn):

did you try ? whats your answer ?

OpenStudy (anonymous):

in all honesty i do not even know how to solve it :/ im just in a rush to finish

OpenStudy (anonymous):

i think its either A or D @hartnn

hartnn (hartnn):

let me give you a hint then \(\Large \sqrt{a^2} = |a|\) we require to put an absolute sign only when the exponent is even number

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

yessss lol thank you :D

hartnn (hartnn):

welcome ^_^